OpenStudy (anonymous):

What are the real and complex solutions of the polynomial equation? x^3-216=0

OpenStudy (anonymous):

a. 3, 3+3i sqrt. 3, 3-3i sqrt. 3 b. 3, -3+3i sqrt. 3, -3-3i sqrt. 3 c. 6, 3i sqrt. 3, -3+3i sqrt. 3 d. 6, -3+3i sqrt. 3, -3-3i sqrt. 3

OpenStudy (anonymous):

fan and medal?

OpenStudy (p0sitr0n):

real x=6 complex x=-3-i3sqrt(3) and x=-3+i3sqrt(3)

OpenStudy (anonymous):

@P0sitr0n

OpenStudy (anonymous):

thank you!